In triangle ABC, the measure of angle A is 141 degrees more than the measure of angle B. The measure of angle B is the same as the measure of angle C. Fine each measure.
Answer by mananth(16949)

let B & C be x
A will be 141+x
x+x+x+141=180
3x+41=180
3x=180-141
3x=39
x=13
13,13, 154